OpenStudy (anonymous):

can someone show me the steps to the answer, I have the answer but need to see it written out

OpenStudy (anonymous):

\[\frac{ 4 }{ x }=\frac{ 9 }{ 5 }-\frac{ 7x-4 }{ 5x }\] with an answer of 8